OpenStudy (anonymous):

Which equation is equivalent to 2x+4a=10 a. x=5/4a b. x=5+2a c. x+10-4a d. x=5-2a

OpenStudy (anonymous):

2x + 4a = 10 2x = 10 - 4a x = 5 - 2a

OpenStudy (anonymous):

thanks u sure

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

yess